OpenStudy (anonymous):

Need Calc Help!!! Please! For each series given, state convergent or divergent. If convergent, find the sum. Justify all answers mathematically. (4^n+1)/(5^n) for n=0 to infinity

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\begin{align*}\sum_{n=0}^\infty \frac{4^n+1}{5^n}&=\sum_{n=0}^\infty \frac{4^n}{5^n}+\sum_{n=0}^\infty \frac{1}{5^n}\\ &=\sum_{n=0}^\infty \left(\frac{4}{5}\right)^n+\sum_{n=0}^\infty \left(\frac{1}{5}\right)^n\end{align*}\] What do you know about geometric series and when they converge/diverge?

OpenStudy (anonymous):

not really general idea but i am not confident with it yet

OpenStudy (sirm3d):

a geometric series \[\Large \sum_{n=0}^\infty r^n\] is convergent if \[\Large -1<r<1\] and divergent otherwise

OpenStudy (sirm3d):

furthermore, if the series is convergent, then \[\Large \sum_{n=0}^{\infty} r^n = \frac{1}{1-r}\]

OpenStudy (anonymous):

Thank you, which series has to do with limits and whether they exist or not